### v1

# Beyond Flat Text: Dual Self-inherited Guidance for Visual Text Generation

Abstract
In real-world images, slanted or curved texts, especially those on cans, banners, or badges, appear as frequently, if not more so, than flat texts due to artistic design or layout constraints. While high-quality visual text generation has become available with the advanced generative capabilities of diffusion models, these models often produce distorted text and inharmonious text background when given slanted or curved text layouts due to training data limitation. In this paper, we introduce a new training-free framework, STGen, which accurately generates visual texts in challenging scenarios (e.g., slanted or curved text layouts) while harmonizing them with the text background. Our framework decomposes the visual text generation process into two branches: (i) Semantic Rectification Branch, which leverages the ability in generating flat but accurate visual texts of the model to guide the generation of challenging scenarios. The generated latent of flat text is abundant in accurate semantic information related both to the text itself and its background. By incorporating this, we rectify the semantic information of the texts and harmonize the integration of the text with its background in complex layouts. (ii) Structure Injection Branch, which reinforces the visual text structure during inference. We incorporate the latent information of the glyph image, rich in glyph structure, as a new condition to further strengthen the text structure. To enhance image harmony, we also apply an effective combination method to merge the priors, providing a solid foundation for generation. Extensive experiments across a variety of visual text layouts demonstrate that our framework achieves superior accuracy and outstanding quality.
Figure 1: STGen for visual text generation in challenging layout. Using a pre-trained visual text generation model (e.g., AnyText [32]), our method, STGen, guides the model to adjust the text region in latent space during image synthesis, producing images that more faithfully represent the input prompt with precise visual text.
11footnotetext: Equal contribution.
1 Introduction
Visual text generation is an emerging yet challenging research area in image generation, because text is fine-grained and difficult to be balanced with image. There are two primary methodologies for visual text generation: local visual text blending and global visual text generation. Local visual text blending methods solely render visual text within given images, represented by early GAN-based methods [3, 33, 22, 12, 31]. Recent SceneVTG [38] can render texts in small regions by utilizing diffusion-based models. In contrast, global visual text generation creates new images with integrated text, raising an inevitable challenge to achieve both accurate visual texts and harmonious image content. Recent studies [5, 28, 1, 15, 18, 34, 16, 32] have attempted to address this challenge by leveraging the capabilities of diffusion models [9, 25, 30, 6, 20, 28, 26, 27, 4]. Among them, AnyText [32] distinguishes itself by producing impressive images integrated with outstanding multilingual text. It opens an era of universal visual text generation using large pretrained Visual Text Generation Model (VTGM). In this paper, we follow in the footsteps of AnyText and focus on global visual text generation.
Although large pretrained VGTMs take a significant step towards universal visual text generation, they still do not possess the capability that is robust to handle variable user input, because of insufficient training data. Their performance is unsatisfactory in challenging cases, such as commonly seen slanted or curved texts in real-world images due to artistic design or layout constraints. Specifically, the subpar performance is reflected in text distortion and background occlusion as shown in Fig. 2.
One naive solution is to train VTGM on a more diversified dataset that covers various text configurations. But it is resource-intensive and time-consuming. To overcome this shortcoming, we propose to use VTGM which is well-trained on simple visual text (e.g. flat visual texts) to guide difficult visual text generation (e.g. slanted or curved visual texts). When plainly using the pretrained VTGM (e.g. AnyText), it can generate a latent that contains clear and readable visual texts even at an early stage given a flat text mask. Nevertheless, in tilted situations, text in the latent becomes blurry and wrong, leading to incorrect results.
Figure 2: Failure cases of AnyText [32]. In the top row, we show examples of 2 failure scenarios: Textual distortion (left) and Background occlusion (right). In the bottom row, we show images obtained when using our method under the same seeds.
To address this issue, we propose a training-free method named Slanted Text Generation (STGen). STGen is a plug-and-play method that corrects text regions in latent space during inference. Specifically, our approach employs a dual-branch framework. The first branch, the Semantic Rectification Branch (SRB), utilizes a latent generated using the same prompt, but with a simplified shape, as a robust semantic prior. This branch simultaneously rectifies distorted text predictions and harmonizes the text with its background. For complex visual text generation, such as text layouts composed of multiple tilted or curved sections, we propose a Divide and Conquer strategy to efficiently reconfigure the text shape and obtain a reasonable semantic prior. The second branch, Structure Injection Branch (SIB), extracts rich structural information from glyphs and injects it into the latent space as a structural prior, further enhancing the accuracy of the visual text. Rather than simply merging the two priors, we adopt a novel combination method for optimized integration. Together, the dual-branch framework offers effective guidance for the generation process. Our method addresses the shortcomings of the existing VTGM, ensuring the generated visual texts are both accurate and well-balanced with the overall image as shown in Fig. 1.
To the best of our knowledge, our method is the first specifically for generating visual texts in complex layouts and achieves the state-of-the-art results without additional training. We conduct thorough experiments to demonstrate the superiority of our method and the effectiveness of each component in generating complex visual texts while enhancing overall image quality and coherence. The main contributions of our work are summarized as follows:
  1. (i)
    We present a new challenge: generating visual text in complex layouts with diffusion models. To tackle this, we introduce STGen, a dual-branch, training-free framework that can be easily integrated with existing text generation models. This framework enables the model to generate text in complex layouts by incorporating both structural and semantic priors.
  2. (ii)
    We introduce an effective method to combine given priors within the latent space. This approach seamlessly merges priors and the latent, ensuring a consistent latent range and resulting in a balanced, high-quality image.
  3. (iii)
    We propose a benchmark extending the AnyText-benchmark to evaluate texts in complex layouts, enabling comprehensive assessment across varying text slant difficulties.
2 Related Works
Denoising diffusion probabilistic models (DDPMs) [9, 25, 30, 6, 20, 28, 26, 27, 4] have made significant contribution to the field of text-to-image synthesis, enabling the generation of high-quality images. Despite advancements in text-to-image synthesis models, achieving precise control over image details remains challenging, particularly in visual text generation.
2.1 Local Visual Text Blending
Local Visual Text Blending focuses on producing localized images that contain a single line of text. Methods in this area often require a background or original image for text generation or editing. SynText [7] detects the region in the image that can put texts on it and generate texts in the corresponding area. Following their work, SceneVTG [38] introduces a model to predict the text generating area and specifically designs a framework that takes line masks and word masks separately to help generate texts in different shapes. SceneVTG is less user-friendly because it requires an erased image and lacks support for multilingual text generation. AnyTrans [21] provides an end-to-end method for local visual text blending, it identifies the texts in the image, erases them from the image, and generates the translated texts in the same area. The above methods require an input image to render or generate, which is good for editing, but cannot generate the whole image from scratch.
2.2 Global Visual Text Generation
Recent studies such as Imagen [28] and e-Diff-I [1] find that simply by replacing the text encoder from CLIP [23] to more capable models like T5 [24] can improve visual text generation. Liu et al. [15] further replace the character-blind text encoder with a character-aware text encoder. Even though replacing text encoders seems to be a straightforward approach, it still struggles to generate complex characters such as Korean, Japanese, and Chinese. GlyphByT5 series [16, 17] employ character-aware ByT5 encoder [14] and a new cross attention mechanism to compute the text region and image region separately. GlyphDraw [18] focuses on generating complex characters such as Chinese and proposes a sound baseline by introducing glyph condition for visual text generation. It uses the glyph image of the text and a location mask as conditions, allowing the model to control both the content and position of the generated text. However, it can generate only one line of text per image. TextDiffuser [5] trains a dedicated layout generation module to specifically generate character-level mask as the condition for the diffusion model, improving performance for Latin texts. But it still cannot generate non-Latin texts. GlyphControl [34] employs ControlNet [36] for visual text generation and uses a rendered glyph image as a control condition, making it possible to enable models like Stable Diffusion to generate texts and maintain their text-to-image capability at the same time. Nonetheless, GlyphControl can only generate texts in straight lines and often produces unwanted texts outside the text region. Following the above works, AnyText [32] proposes a robust and unified framework for visual text generation. It can generate high-quality multilingual texts in images, thanks to the text perceptual loss which takes the correctness of the generated text into account during training. However, AnyText performs poorly for generated text leaning more than $45$ degrees. TextGen [35] analyzes how the control information influences the generation in different timesteps. Diff-Text [37] uses off-the-shelf diffusion model and canny controlnet to generate visual texts, where the canny controlnet utilizes the canny image of text glyph to control visual text generation. Nonetheless, Diff-Text binds text position to predefined objects like ’sign’ or ’billboard’ in the prompt. It may fail if a prompt does not include predefined objects, leading to unnatural images.
3 Method
Figure 3: Pipeline of our method. Given user input on the leftmost side, which contains a prompt and a mask $l_{p}$ specifying positions for generating visual texts, we first split the $l_{p}$ using Divide and Conquer Strategy and obtain glyph image $l_{g}$ and flat position mask $\tilde{l}_{p}$. Then $\tilde{l}_{p}$ and $l_{g}$ are input to Semantic Rectification Branch and Structure Injection Branch respectively. In Semantic Rectification Branch, based on $\tilde{l}_{p}$, we render flat glyph $\tilde{l}_{g}$ and use them along with prompt and random noise $z_{T}$ to generate the latent with flat visual text. This latent serves as a semantic prior, providing rich semantic information for both the generation of the text and its background. $l_{g}$, on the other hand, is converted into the latent space as a structural prior for structural refinement of the text. Finally, the two prior combined together to guide the generation of the visual text in $l_{p}$.
Given a prompt $y$ describing the image and a challenging position mask $l_{p}$, our goal is to generate an image based on $y$ that incorporates visual text at the specified positions dictated by $l_{p}$. Current visual text generation models struggle to generate visual texts in more challenging $l_{p}$ settings, such as tilted or curved texts layouts.
Our method aims to solve the problem based on two key insights: First, when the position mask $l_{p}$ is flat, current models can generate visual texts with high accuracy. We leverage this proficiency with flat visual texts to address challenging scenarios, providing a strong semantic prior for generation. Second, the glyph image contains little semantic information but is rich in glyph structural details. It can serve as a structure prior to further refine the structural information of the visual texts within the latent, thus enabling it to generate accurate slanted visual texts.
Our approach consists of two main branches: the Semantic Rectification Branch and the Structure Injection Branch. As shown in Fig. 3, the semantic rectification branch first takes the masks $\tilde{l}_{p}$ reconfigured from $l_{p}$, prompt $y$, and random noise $z_{T}$ to generate the latent representation $z_{0}^{f}$, which contains reconfigured and flat visual text as a semantic prior. The structure injection branch then takes the rendered glyph image $l_{g}$ and $z_{t}$ to generate a structural prior. These two priors are merged and fed into the VTGM, along with $y$, $l_{p}$, and $l_{g}$, for the denoising process. Further details of these branches are discussed in Sec. 3.1 and Sec. 3.2.
3.1 Semantic Rectification Branch
In the initial stages of the DDIM denoising process in the VGTM, clear text is created when applying a flat position mask, while blurry and unclear text is generated when using a slanted mask. The reason for the distortion lies in the semantic drift in the text region when tilted. Motivated by this observation, we inherit the high quality flat text generation capabilities of the existing visual text model, adopting the latent of flat text as a constant reference to rectify the semantic information for challenging text generation.
Reference Branch for Semantic Rectification.
Parallel to the generation branch, a separate branch is employed to generate flat visual texts using the same prompt $y$ for additional semantic information, which we refer to as the semantic rectification branch. In this branch, the flat visual text latent is blended into the tilted one within the text region, while the rest of the latent remains unchanged.
As shown in Fig. 3, the flat position $\tilde{l}_{p}$ and corresponding glyph image $\tilde{l}_{g}$ are fed into the VTGM along with prompt $y$ and random noise $z_{T}$. After $T$ denoising steps, we obtain the reference latent $z^{f}_{0}$. We rotate $z^{f}_{0}$ to match the user-given position, and extract its text region $z_{f}$ as a semantic guide for the branch below:
$\tilde{z}_{t}=z_{f}\odot l_{p}+z_{t}\odot(1-l_{p}),$ (1)
where $l_{p}$ is the position mask input by user. This operation effectively rectifies the visual text using the accurate text in the $z_{f}$. Thanks to the faithful background semantic information embedded in the $z_{f}$, we significantly reduce the erroneous non-textual semantic information in the text region while maintaining a coherent background, thereby avoiding background occlusion.
AdaIN Combination.
AdaIN [10] is originally developed for style transfer tasks. It substitutes mean and standard deviation of the source feature with those of the target feature. Masui et al. [19] demonstrated that AdaIN can be applied directly to Diffusion models for style transfer without any additional training. In our method, we also incorporate AdaIN, but with the goal of preserving image integrity against disruptions in latent distribution caused by the replacement operation:
$\tilde{z}_{t}$ $={\rm AdaIN}(z_{f},z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$ (2)
$=(\sigma(z_{t})(\frac{x-\mu(z_{f})}{\sigma(z_{f})})+\mu(z_{t}))\odot l_{p}+z_{t}\odot(1-l_{p}),$
where $\mu(\cdot)$ and $\sigma(\cdot)$ denote the channel-wise mean and standard deviation. The aid of AdaIN ensures the latent for guidance has the same range as the original latent, minimizing the change in latent distribution. With a more consistent latent, the visual text semantic information is further enhanced while preserving the overall image coherence.
Divide and Conquer Strategy.
In practice, complex position masks are often provided for generating realistic visual texts, which often consist of multiple straight or curved parts, bringing obstacles to the geometric alignment to flat position masks $\tilde{l}_{p}$ in our method. To deal with the problem, we propose a divide-and-conquer strategy to effectively segment complex position masks into more manageable straight sections as shown in Fig. 3. We use Bézier curves to define the upper and lower boundaries of $l_{p}$ and establish a baseline for the texts by averaging these curves. We identify splitting points on the curve where the direction vectors at each point are parallel to the boundaries of the minimum bounding box of $l_{p}$, therefore creating $N$ reconfigured masks:
$l_{p}=\left\{l^{0}_{p},l^{1}_{p},\ldots,l^{N-1}_{p}\right\}.$ (3)
Subsequently, we rotate and regroup these reconfigured masks to obtain flat position masks $\tilde{l}_{p}$ and render corresponding glyphs $\tilde{l}_{g}$:
$\tilde{l}_{p}=\left\{\tilde{l}^{0}_{p},\tilde{l}^{1}_{p},\ldots,\tilde{l}^{N-1}_{p}\right\},\tilde{l}_{g}=\left\{\tilde{l}^{0}_{g},\tilde{l}^{1}_{g},\ldots,\tilde{l}^{N-1}_{g}\right\}.$ (4)
With the assistance of this branch, we can generate highly accurate texts and create harmonious, diverse images by leveraging rich semantic information in both the visual texts and background.
1 Input: A text prompt $y$, an input position image $l_{p}$, a flat reference position image $\tilde{l}_{p}$ and a set of iterations for refinement $\left\{t_{1},\ldots,t_{k}\right\}$ and a trained Visual Text Generation Model $\mathcal{M}$
2 Output: Denoised vector $z_{0}$
3   Render the glyph image $l_{g}$ and compute its latent $z_{g}$;
4 Compute the reference latent $z^{f}_{0}$;
5 Rotate and crop $z^{f}_{0}$ to get $z_{f}$ according to the transformation from $\tilde{l}_{p}$ to $l_{p}$;
6 for $t=T,T-1,\ldots,1$ do
7       if $t\in\left\{t_{1},\ldots,t_{k}\right\}$ then
8             $\tilde{z}_{t}\leftarrow{\rm AdaIN}(z_{f},z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$;
9             $\hat{z}_{t}\leftarrow\rho{\rm AdaIN}(z_{g},z_{t})+(1-\rho)\tilde{z}_{t}$;
10             $\ddot{z}_{t}\leftarrow(\kappa_{t}\lambda\hat{z}_{t}+(1-\kappa_{t})z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$;
11             Set $z_{t}\leftarrow\ddot{z}_{t}$;
12
13       end if
14      Set $z_{t-1}\leftarrow\mathcal{M}(z_{t},y,t,l_{p},l_{g})$;
15
16 end for
17
18Return $z_{0}$
Algorithm 1 STGen
3.2 Structure Injection Branch
Despite semantic guidance, the predictions may still deviate due to limited information in structure, causing cumulative errors and structural inconsistencies that require enriched structural guidance for clarity. To resolve this, we introduce a glyph structure prior directly into the latent space. Unlike early methods [18] that use glyph images directly as auxiliary information, we consider text glyphs as essential parts of the image, which should be comprehended by the visual text generation model in the latent space. As shown in Fig. 3, the structure injection branch feeds the rendered glyph image $l_{g}$ to Variational Autoencoder (VAE) [11] encoder to obtain the latent $z_{g}$. As an extension of semantic prior, we incorporate the $z_{g}$ as a structure prior into the latent $z_{t}$, using the same AdaIN operation above to regulate the range, which is represented as:
$\hat{z}_{t}={\rm AdaIN}(z_{g},z_{t})\odot l_{p}+z_{t}\odot(1-l_{p}).$ (5)
This operation provides the model with a structural foundation that serves as a strong starting point in the generation process, further enhancing the overall structure of the visual texts.
Combining the two branches, the merged prior is represented by:
$\hat{z}_{t}=\rho{\rm AdaIN}(z_{g},z_{t})+(1-\rho)\tilde{z}_{t},$ (6)
where $\rho$ is a hyper-parameter that adjusts the balance between semantic and structural information. The modified latent is represented as follows:
$\ddot{z}_{t}=(\kappa_{t}\lambda\hat{z}_{t}+(1-\kappa_{t})z_{t})\odot l_{p}+z_{t}\odot(1-l_{p}),$ (7)
where $\lambda$ is a hyper-parameter and the $\kappa_{t}$ is a temporal factor that decays over time with each timestep. Together, they control the injection strength of the merged prior. The complete algorithm is represented as Algorithm 1.
4 Experiments
[…]
English -0.5 0.5 44.88 64.11 0.3005
0.5 0.5 45.43 65.10 0.3027
0.5 0.25 45.12 63.85 0.3036
0.5 0.75 45.10 65.65 0.3006
0.5 1.50 45.01 65.34 0.3009
0.5 2.00 44.53 65.17 0.3009
Chinese -0.5 0.5 49.28 87.29 0.3067
0.5 0.5 49.96 88.08 0.3071
0.5 0.25 50.70 87.86 0.3076
0.5 0.75 49.05 87.89 0.3058
0.5 1.50 47.40 87.64 0.3061
0.5 2.00 47.88 87.63 0.3059
Table 1: Sensitivity analysis for $[\lambda,\rho]$.
[…]
easy medium hard total easy medium hard total
English TextDiffuser [5] 49.88 20.10 0.262 29.06 70.86 41.25 4.47 45.78 0.3091
GlyphControl [34] 19.00 1.63 0.22 9.47 41.74 9.17 2.98 22.95 0.3206
SD1.5+SceneVTG [38] 13.62 4.97 1.12 8.06 24.14 13.40 3.00 15.80 0.3112
Anytext [32] 62.77 29.12 2.02 38.04 83.64 56.99 14.23 58.59 0.3007
Diff-Text [37] 40.13 26.42 8.85 28.38 61.44 45.37 17.94 45.91 0.2962
GlyphControl+Ours 39.19 2.63 1.05 19.48 55.57 11.83 6.74 31.17 0.3175
AnyText+Ours 71.25 37.25 6.60 45.43 87.54 64.42 24.56 65.10 0.3027
Chinese TextDiffuser [5] 5.41 4.07 0.09 4.07 56.81 49.37 43.01 52.47 0.3066
GlyphControl [34] 2.33 0.24 0.0 1.41 50.98 15.58 35.94 40.24 0.3192
SD1.5+SceneVTG [38] 1.60 0.81 0.18 1.14 7.33 4.62 4.61 6.20 0.3025
Anytext [32] 66.00 26.38 2.02 44.78 94.48 81.37 59.38 84.74 0.3064
Diff-Text [37] 23.46 17.26 7.35 18.95 68.27 62.71 48.65 63.21 0.2950
GlyphControl+Ours 6.27 0.57 0.74 3.93 62.50 18.08 45.58 49.40 0.3168
AnyText+Ours 69.22 35.58 8.55 49.96 95.37 85.85 68.77 88.08 0.3071
Table 2: Quantitative Comparison between STGen and other competitors on both English and Chinese sets. All competitors are evaluated based on their officially released models. Numbers in bold indicate the best performance, and underscored numbers indicate the second best.
Baseline Baseline Preference Our Method Preference
TextDiffuser 23.07% 76.93%
GlyphControl 21.96% 78.04%
SD1.5+SceneVTG 18.99% 81.01%
AnyText 25.31% 74.69%
Diff-Text 17.00% 83.00%
Table 3: User study results. The participants were asked to select the best results based on image quality, the accuracy of the generated visual texts, and the similarity between the prompt and the generated images.
4.1 Implementation Details
Our framework is based on the pretrained AnyText model. During inference, we utilize a single RTX 3090 GPU for images in size of (512, 512). The framework achieves optimal performance when $\lambda$ is within the range $[-0.5,0.5]$ and $\rho$ is within $(0,2]$. We set $\kappa_{t}=10^{t-T}$ for effective guidance and harmonious text region boundary. Specifically, when $\lambda$ is in $[-0.5,0)$, the diversity of generated visual texts increases, while $\lambda$ in $(0,0.5]$ increases the variety of the image background. As shown in Tab. 1, our method demonstrates robustness with respect to hyper-parameter sensitivity. In subsequent experiments, we set $\lambda$ to $0.5$ and $\rho$ to $0.5$ which has a balanced performance on both Chinese and English for evaluation.
4.2 Evaluation Setup
Since there are currently no openly available datasets that focus on challenging visual texts generation, we establish a new benchmark for evaluation based on the AnyText-benchmark [32]. For each text position mask, we randomly rotate the masks from the original benchmark and resolve any intersections, resulting in 984 prompts for LAION-word (English evaluation) and 919 prompts for Wukong-word (Chinese evaluation), each with corresponding position masks. Each prompt includes at least one visual text line and up to five visual text lines. During evaluation, we categorize the masks into three levels based on their angles: ‘easy’ (0°–30°), ‘medium’ (30°–60°), and ‘hard’ (60°–90°), facilitating multi-level assessments.
Textual accuracy and background-text coherence are two main factors that determine the quality of slanted text generation, which we quantitatively evaluate through OCR accuracy. Following AnyText [32], we select the following two metrics for comparing OCR accuracy at word-level and character-level, respectively: (1) Sentence Accuracy (Sen.Acc); (2) Normalized Edit Distance (NED).
We evaluated existing competing methods, including TextDiffuser [5], GlyphControl [34], SceneVTG [38], AnyText [32] and Diff-Text [37] using the benchmark and metrics mentioned above. It is worth noticing that the SceneVTG [38] cannot generate image from scratch. Therefore, we use the images generated by Stable Diffusion [27] with the same prompt as background image. For all the methods in the evaluation, we use the preprocessed mask using the Divide and Conquer strategy mentioned above to generate images to ensure fairness.
Figure 4: Qualitative comparison of our method and state of the art models in both English and Chinese text generation.
4.3 Qualitative Comparisons
As presented in Fig. 4, TextDiffuser tends to generate low-quality images, specifically in the second and the fifth columns. GlyphControl often produces texts outside the given text position, as can be seen in the fourth column. SceneVTG [38] exhibits severe text distortion, particularly in the first and third columns where the texts are extremely distorted and unrecognizable. Typical background occlusion is also observed in the second, third, and fourth columns, where the texts “Burning”, “Mamba” and “Twinkle” are obscured by non-textual elements such as lava, a human figure, and stars, respectively. This indicates that SceneVTG can only render texts against simple and clear backgrounds and does not integrate well with other text-to-image models. For Diff-text [37], there is a noticeable lack of coherence in the images due to insufficient consideration for harmony between the visual texts and the background, as particularly evidenced in the last 2 columns where the texts are rendered outside the bow and the watch. In AnyText [32], the generated visual texts tend to be distorted, this was caused by the loss of semantic and structure information during the inference. Moreover, in the first column, the text “Track” occludes with the train head, affecting both the text and the train. For our method, we can generate images with superior quality with accurate English and Chinese visual texts integrated even in complex scenarios like in the fourth and fifth columns. It’s worth noting that except for resolving text distortion, our method also improves the background occlusion between visual text and non-textual elements. For instance, in the third column of Fig. 4, the text “Mamba” is originally disrupted by the human head using AnyText [32], while our method pushes the human head aside to generate a clear text and reasonable non-textual element.
4.4 Quantitative Analysis
OCR Accuracy.
As shown in Tab. 2, our method shows substantial improvements over the competitive methods in both English and Chinese across all levels. Notably, our method doesn’t require additional training and enhances the performance of GlyphControl and AnyText at the easy level. At the hard level, AnyText with our method yields results closest to the ground truth in the NED metric, with an improvement of approximately $10\%.$ over the baseline AnyText in both languages.
Text-Image Similarities.
In the absence of ground-truth images, we use the CLIP score [8] to assess the consistency between the prompt and the generated image. We compute the average cosine similarity between the prompt and the generated image, excluding the influence of the visual texts. As shown in Tab. 2, our method increases the accuracy of visual text generation without compromising the baseline performance. Additionally, our method improves the text-image similarity slightly, due to our design that minimizes conflicts between the image composition and the visual text. This leads to a more coherent layout for both the image and the text.
User Study.
Following [2], we conduct user study to comprehensively compare the generation results in Tab. 3. We create 27 input sets with varying levels of difficulty in visual text generation. For each input set, we provide the respondents with our input prompt and position mask and two image alternatives: our result and a baseline (in random order). We ask them to select which image has the best quality and has the most accurate visual text. The final score is calculated based on the times the respondent choose the approach averaged on the total prompts. We collect 126 user judgments from a diverse group composed of both experts in the field and individuals with no specific background and report the percentage of votes. As shown in Tab. 3, our method is the most favored among all the circumstances.
5 Ablation Study
Figure 5: Ablation Study Visualization. The first column shows accuracy improvement with SRB. The second column demonstrates that while SIB enhances accuracy, it compromises coherence. The third column highlights that combining both branches yields accurate, coherent visual texts. Finally, AdaIN combination in the fourth column further refines image quality.
In this section, we demonstrate the effectiveness of our components through an ablation study on the medium level. We use OCR accuracy (Sen.Acc and NED) and CLIP score as the main metrics, given their importance in evaluating visual text generation. All parameters are kept consistent with those outlined in Sec. 4.1. The analysis of each component will be presented in the following parts of this section.
SRB without AdaIN.
As shown in the first two rows of Tab. 4, the addition of Semantic Rectification Branch improves the visual text accuracy. Similarly, as demonstrated by the first column of Fig. 5, the generated visual text has a clear structure. However, this also negatively impacts text-image consistency, as indicated by the slight drop in the CLIP score. This occurs because the operation of replacing partial latent disrupts the latent distribution and affects the representation of other parts of the image. Thanks to the rich semantic information in the latent, the drop is minor.
SRB SIB AdaIN Sen. Acc NED CLIP Score
29.12 56.99 0.3007
36.22 61.17 0.3003
36.93 64.35 0.3006
37.25 64.42 0.3027
Table 4: Ablation study. The results validate the effectiveness of each component.
SIB without AdaIN.
As illustrated in the second and third rows of Tab. 4, Structure Injection Branch further improves the accuracy. As similarly shown in the third column of Fig. 5, this branch can structurally improve the text structure, resulting in impressive improvement. However, when applied alone, it may disrupt image coherence, as shown in the second column and third row of Fig. 5, where the word ’Fun’ occludes the eye. This issue arises from the lack of semantic information in the structural prior.
AdaIN Combination.
As revealed by the third and fourth column of Fig. 5, AdaIN improves the accuracy and the harmony of the text and its background, lifting the CLIP score to a new level in the Tab. 4.
6 Conclusion
In this work, we push the boundary of visual text generation by tackling the challenge of creating complex visual texts. We present a dual-branch method STGen, which decomposes the visual text generation into 2 branches: Semantic and Structure. In Semantic Rectification Branch, our method leverages the rich semantic information composed in the latent when generating accurate visual texts in simple scenarios to guide the generation in complex scenarios. In the Structure Injection Branch, we polish the visual text structure using the structural information extracted from glyphs at the same time. For challenging scenarios, we break them down into simpler steps to manage effectively. This approach can be integrated seamlessly with the existing visual text generation model using a dedicated ControlNet for generation. To exhibit its superior performance, we conduct extensive experiments on our proposed benchmark, demonstrating its excellence. Our approach offers a straightforward yet effective way to enhance the capabilities of existing visual text generation models. We believe STGen sets the stage for real-world applications, enabling more accurate and sophisticated visual text generation.
Supplementary Material
Figure 6: Comparison of predicted $x_{0}$ under different inference steps in various masks. The first column is the AnyText intermediate results under a flat mask, AnyText can predict the structure of the texts at the very early stage of inference. The rest of the columns show intermediates in a slanted mask using AnyText and other operations using different methods. Our method can revise the originally wrong texts and harmonize the background.
7 Additional Background
Latent Diffusion Models.
The Latent Diffusion Models (LDM) denoise a noisy vector $z_{t}$ to $z_{0}$ that is mapped to an image $x_{0}$ through an autoencoder, based on a text prompt $y$. To sequentially remove the noise $\varepsilon$, a network $\varepsilon_{\theta}$ is trained to minimize the loss:
$\mathcal{L}=\mathbb{E}_{z_{0},y,\varepsilon\sim\mathcal{N}(0,1),t}||\varepsilon-\varepsilon_{\theta}(z_{t},t,c(y))||_{2}^{2},$ (8)
where the $c(y)$ is the conditioning embedding of the prompt $y$, $z_{t}$ is a noisy vector obtained by adding noise to $z_{0}$ according to the timestep $t$. During inference, provided a random noise vector $z_{T}$, the trained network iteratively removes the predicted noise to produce a latent $z_{0}$ for $T$ steps. Namely, we employ popular DDIM sampling [29] for inference:
$z_{t-1}=$ $\sqrt{\alpha_{t-1}}\underbrace{(\frac{z_{t}-\sqrt{1-\alpha_{t}\epsilon_{\theta}(z_{t})}}{\sqrt{\alpha_{t}}})}_{\text{predicted $z_{0}$}}$ (9)
$+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}\cdot\epsilon_{\theta}(z_{t})}.$
The DDIM Sampling in general combines the predicted $z_{0}$ and $z_{t}$ to obtain the $z_{t-1}$. We mainly analyze the predicted $x_{0}$ that is decoded from predicted $z_{0}$ in the rest of the section.
Adaptive Instance Normalization.
The Adaptive Instance Normalization (AdaIN) [10] is previously adopted in the task of style transfer. It substitutes mean and standard deviation of the activations from each CNN [13] filter of the original image with those of the style image as follows:
${\rm AdaIN}(x,y)=\sigma(y)(\frac{x-\mu(x)}{\sigma(x)})+\mu(y),$ (10)
where $x$ and $y$ stand for the activation of a CNN filter from the content image and style image, $\mu(\cdot)$ and $\sigma(\cdot)$ are channel-wise mean and standard deviation. Masui et al. [19] further prove that the AdaIN can be directly applied to the LDM without any additional training for style transfer.
8 Analysis of Predicted $x_{0}$
As shown in Fig. 6, AnyText can produce clear and intact visual texts in structure even in the first inference step. However, once tilted, the prediction becomes blurry and wrong as in the second column. The following columns show the effectiveness of our method. SRB can improve the structure. The SIB further improves the structure but produces visual texts that are not coherent with the background.
Figure 7: Results under different lambda. When $\lambda$ is -0.5, we obtain colorful and styled visual texts in the first column. When $\lambda$ is 0.5, we produce images with rich background. When $\lambda$ is 0, the method degrades to vanilla AnyText.
Language $\lambda$ $\rho$ Sen.ACC NED CLIP score
easy medium hard total easy medium hard total
English -0.5 0.5 71.33 35.51 6.14 44.88 87.23 62.98 22.79 64.11 0.3005
0.5 0.5 71.25 37.25 6.60 45.43 87.54 64.42 24.56 65.10 0.3027
0.5 0.25 71.37 37.11 5.32 45.12 87.10 63.80 21.19 63.85 0.3036
0.5 0.75 69.60 38.39 7.16 45.10 87.33 65.54 25.94 65.65 0.3006
0.5 1.25 61.48 35.16 6.33 40.18 85.38 63.73 25.44 64.11 0.3008
0.5 1.5 69.47 38.92 6.48 45.01 87.23 65.03 25.43 65.34 0.3009
0.5 1.75 69.19 38.17 6.56 44.70 87.18 65.31 25.40 65.39 0.3011
0.5 2.0 68.88 37.96 6.71 44.53 87.07 64.85 25.26 65.17 0.3009
Chinese -0.5 0.5 68.61 35.02 7.54 49.28 95.17 85.23 66.05 87.29 0.3067
0.5 0.5 69.22 35.58 8.55 49.96 95.37 85.85 68.77 88.08 0.3071
0.5 0.25 71.35 35.42 6.16 50.70 95.46 85.88 67.37 87.86 0.3076
0.5 0.75 66.12 37.79 10.66 49.05 94.54 85.77 70.38 87.89 0.3058
0.5 1.25 64.28 38.11 10.57 48.03 94.31 86.14 71.58 88.07 0.3060
0.5 1.5 63.85 37.05 9.83 47.40 94.24 85.52 70.26 87.64 0.3061
0.5 1.75 65.53 38.11 10.29 48.12 94.29 85.85 70.21 87.73 0.3064
0.5 2.0 65.02 35.91 10.11 47.88 94.44 85.37 69.78 87.63 0.3059
Table 5: Detailed sensitivity analysis for $[\lambda,\rho]$.
9 Additional Results
Here we provide additional results of our method. Our code will be released upon acceptance.
9.1 Additional Quantitative Results
Hyper-parameters Sensitivity
As shown in the Tab. 5, our method shows robustness in hyper-parameters change. The best performance is obtained when $\rho$ is 0.75 at the hard level. But the method achieves the best performance in the easy level of Chinese and English set when $\rho$ is 0.25. To ensure fairness, we choose the performance when $\rho$ is 0.5 which has a balanced performance for comparison.
9.2 Additional Qualitative Results
Choice of $\lambda$
We can observe in Fig. 7 that when $\lambda$ is in the range of $(0,0.5]$ we can achieve background with more diversity and in the range of $[-0.5,0)$ the diversity of texts is increased consistent with what we mentioned in Sec. 4.1.
Coherence with Background
As shown in Fig. 8, thanks to the rich semantic information contained in the latent with flat visual texts, we can generate visual texts that are perfectly blended in with the background. The text ‘Harvest’ in the bottom right is more harmonious with the background and is closest to the description “interwoven”.
Visual Text Diversity
As shown in the first row of Fig. 10, with proper control, our method can generate texts in various styles and colors. The second row shows that the generated visual texts remain to be accurate under 2 different image styles. The third row demonstrates the robustness of our method under different challenging masks. The fourth row shows our ability to generate various visual texts using the same prompt.
Additional Comparison with baseline.
We present more qualitative comparisons on simple slanted or curved situations in Fig. 11 and Fig. 12, and complex scenarios in Fig. 13. Our method can generate high-quality images with accurate visual texts in all the scenarios.
10 Details of User Study
An illustration of our user study is provided in Fig. 9. Since the caption in the English (LAION) set are often hard for users to understand (e.g. “carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel”). To produce meaningful prompts that can help the user understand, we employ LLMs to generate image prompts for user study. Here we provide a simplified template of our instructions and examples:
1.Task instruction
I am working on a text-to-image generation task. Each Prompt should include a scene description and contain one or two sets of words that need to be present in the image, enclosed in double quotes.
2. Examples
A raccoon stands in front of the blackboard with the words “Deep Learning” written on it …… 3.Trigger CoT reasoning ability of LLMs
Reasoning: Let’s think step by step……
11 Limitations
Our method still inherits the problems of the baseline AnyText, such as the unsatisfactory ability to generate texts on a small scale or to generate texts in a specific font. Restricted by the latent size, our method may not acquire accurate texts when the text region is too small.
Figure 8: Examples of texts blend perfectly with background. Visual text generated by AnyText in the first row is often blurry and wrong. While our method can not only generate accurate visual text but the visual text is also blended with the background perfectly.
Figure 9: User study print screen.
Figure 10: Diversity demonstration.The first row demonstrates various fonts and colors a single prompt can generate. The second row shows same texts with different prompts can still generate coherent and reasonable images. We use different masks in the third row. The fourth row illustrates the ability to simply change the language of the generated texts.
Figure 11: Additional slanted results for our method.
Figure 12: Additional curved results for our method.
Figure 13: Additional complex results for our method.

### v2

@@
 In real-world images, slanted or curved texts, especially those on cans, banners, or badges, appear as frequently, if not more so, than flat texts due to artistic design or layout constraints.
-While high-quality visual text generation has become available with the advanced generative capabilities of diffusion models, these models often produce distorted text and inharmonious text background when given slanted or curved text layouts due to training data limitation.
-In this paper, we introduce a new training-free framework, STGen, which accurately generates visual texts in challenging scenarios (e.g., slanted or curved text layouts) while harmonizing them with the text background.
+While high-quality visual text generation has become available with the advanced generative capabilities of diffusion models, these models often produce distorted text and inharmonious text backgrounds when given slanted or curved text layouts due to training data limitations.
+In this paper, we introduce a new framework, STGen, which accurately generates visual texts in challenging scenarios (e.g., slanted or curved text layouts) while harmonizing them with the text background.
 Our framework decomposes the visual text generation process into two branches: (i) Semantic Rectification Branch, which leverages the ability in generating flat but accurate visual texts of the model to guide the generation of challenging scenarios.
@@
 Extensive experiments across a variety of visual text layouts demonstrate that our framework achieves superior accuracy and outstanding quality.
+The code will be available at https://github.com/Tony-Lowe/STGen.
 Figure 1: STGen for visual text generation in challenging layout.
@@
 1 Introduction
-Visual text generation is an emerging yet challenging research area in image generation, because text is fine-grained and difficult to be balanced with image.
-There are two primary methodologies for visual text generation: local visual text blending and global visual text generation.
-Local visual text blending methods solely render visual text within given images, represented by early GAN-based methods [3, 33, 22, 12, 31].
-Recent SceneVTG [38] can render texts in small regions by utilizing diffusion-based models.
-In contrast, global visual text generation creates new images with integrated text, raising an inevitable challenge to achieve both accurate visual texts and harmonious image content.
-Recent studies [5, 28, 1, 15, 18, 34, 16, 32] have attempted to address this challenge by leveraging the capabilities of diffusion models [9, 25, 30, 6, 20, 28, 26, 27, 4].
-Among them, AnyText [32] distinguishes itself by producing impressive images integrated with outstanding multilingual text.
+Visual text generation is an emerging yet challenging research area in image generation because text is fine-grained and difficult to balance with the image.
+The current methods [3, 37, 25, 15, 35, 43, 6, 31, 1, 18, 21, 38, 19, 36, 5] proposed a series techniques to address it.
+Among them, diffusion based methods (AnyText [36], GlyphControl [38], DiffText [41] etc.) can create new images with integrated text, raising an inevitable challenge to achieve both accurate visual texts and harmonious image content.
+Notably, AnyText [36] distinguishes itself by producing impressive images integrated with outstanding multilingual text.
 It opens an era of universal visual text generation using large pretrained Visual Text Generation Model (VTGM).
-In this paper, we follow in the footsteps of AnyText and focus on global visual text generation.
-Although large pretrained VGTMs take a significant step towards universal visual text generation, they still do not possess the capability that is robust to handle variable user input, because of insufficient training data.
-Their performance is unsatisfactory in challenging cases, such as commonly seen slanted or curved texts in real-world images due to artistic design or layout constraints.
-Specifically, the subpar performance is reflected in text distortion and background occlusion as shown in Fig.
+Although large pretrained VGTMs take a significant step towards universal visual text generation, they still struggle to handle users’ diverse inputs, such as commonly seen slanted or curved texts in real-world images.
+Given these user inputs, the model often leads to text distortion and background occlusion, as shown in Fig.
 2.
+This is because the latent space gradually becomes blurry and distorted (as shown in Fig.
+3) due to insufficient data in such scenarios, thus the model cannot effectively maintain the structure and semantic information in the text region as it does when processing a flat mask.
 One naive solution is to train VTGM on a more diversified dataset that covers various text configurations.
-But it is resource-intensive and time-consuming.
-To overcome this shortcoming, we propose to use VTGM which is well-trained on simple visual text (e.g.
-flat visual texts) to guide difficult visual text generation (e.g.
-slanted or curved visual texts).
-When plainly using the pretrained VTGM (e.g.
-AnyText), it can generate a latent that contains clear and readable visual texts even at an early stage given a flat text mask.
-Nevertheless, in tilted situations, text in the latent becomes blurry and wrong, leading to incorrect results.
-Figure 2: Failure cases of AnyText [32].
-In the top row, we show examples of 2 failure scenarios: Textual distortion (left) and Background occlusion (right).
-In the bottom row, we show images obtained when using our method under the same seeds.
-To address this issue, we propose a training-free method named Slanted Text Generation (STGen).
-STGen is a plug-and-play method that corrects text regions in latent space during inference.
+But it is resource-intensive and data distribution across multiple text layouts cannot be guaranteed.
+Even if they are trained on those data, the results may not be satisfying [32, 10].
+Figure 2: Failure cases of AnyText [36]. The top row illustrates two failure cases: textual distortion (left) and background occlusion (right). The bottom row displays results using our method.
+To overcome this shortcoming, we propose a plug-and-play method named Slanted Text Generation (STGen), which corrects text regions in latent space during inference.
 Specifically, our approach employs a dual-branch framework.
@@
 The second branch, Structure Injection Branch (SIB), extracts rich structural information from glyphs and injects it into the latent space as a structural prior, further enhancing the accuracy of the visual text.
-Rather than simply merging the two priors, we adopt a novel combination method for optimized integration.
-Together, the dual-branch framework offers effective guidance for the generation process.
-Our method addresses the shortcomings of the existing VTGM, ensuring the generated visual texts are both accurate and well-balanced with the overall image as shown in Fig.
-1.
-To the best of our knowledge, our method is the first specifically for generating visual texts in complex layouts and achieves the state-of-the-art results without additional training.
+Rather than simply merging the two priors, we adopt a novel combination method for optimized integration for better coherence in latent space.
+Together, the dual-branch framework offers effective guidance for the generation process without requiring additional training.
+To the best of our knowledge, our method is the first tailored specifically for generating visual texts in complex layouts and achieves state-of-the-art results as shown in Figs.
+1, 5 and 6 without additional training.
+Our framework effectively improves visual text accuracy across various approaches, including early-stage GlyphControl, DiffText, which receives no specified training for visual text generation, and AnyText.
 We conduct thorough experiments to demonstrate the superiority of our method and the effectiveness of each component in generating complex visual texts while enhancing overall image quality and coherence.
@@
     We propose a benchmark extending the AnyText-benchmark to evaluate texts in complex layouts, enabling comprehensive assessment across varying text slant difficulties.
+Figure 3: Comparison of predicted $x_{0}$ under different inference steps.
+The first and second rows show AnyText’s intermediate predicted $x_{0}$ for flat and slanted masks, respectively.
+While predictions remain stable under flat masks, the $x_{0}$ prediction drifts during inference with slanted masks.
+Our method effectively guides the model to maintain accuracy in visual texts.
 2 Related Works
-Denoising diffusion probabilistic models (DDPMs) [9, 25, 30, 6, 20, 28, 26, 27, 4] have made significant contribution to the field of text-to-image synthesis, enabling the generation of high-quality images.
-Despite advancements in text-to-image synthesis models, achieving precise control over image details remains challenging, particularly in visual text generation.
+Denoising diffusion probabilistic models (DDPMs) [12, 28, 34, 8, 23, 31, 29, 30, 4] have made significant contribution to the field of text-to-image synthesis.
+However, achieving precise control over image details remains challenging, particularly in visual text generation.
 2.1 Local Visual Text Blending
-Local Visual Text Blending focuses on producing localized images that contain a single line of text.
-Methods in this area often require a background or original image for text generation or editing.
-SynText [7] detects the region in the image that can put texts on it and generate texts in the corresponding area.
-Following their work, SceneVTG [38] introduces a model to predict the text generating area and specifically designs a framework that takes line masks and word masks separately to help generate texts in different shapes.
+Local Visual Text Blending methods synthesize localized images with single-line textual content.
+Existing approaches typically necessitate a background image for text insertion or modification.
+SynText [9] identifies image regions suitable for text placement and renders text accordingly.
+Following their work, SceneVTG [43] introduces a model to predict the text-generating area and specifically designs a framework that takes line masks and word masks separately to help generate texts in different shapes.
 SceneVTG is less user-friendly because it requires an erased image and lacks support for multilingual text generation.
-AnyTrans [21] provides an end-to-end method for local visual text blending, it identifies the texts in the image, erases them from the image, and generates the translated texts in the same area.
+AnyTrans [24] detects, removes, and replaces text with translations in the same region.
 The above methods require an input image to render or generate, which is good for editing, but cannot generate the whole image from scratch.
 2.2 Global Visual Text Generation
-Recent studies such as Imagen [28] and e-Diff-I [1] find that simply by replacing the text encoder from CLIP [23] to more capable models like T5 [24] can improve visual text generation.
+Recent studies, such as Imagen [31], demonstrate that replacing CLIP text encoder [26] with more advanced models like T5 [27] enhances visual text generation.
 Liu et al.
 [18] further replace the character-blind text encoder with a character-aware text encoder.
-Even though replacing text encoders seems to be a straightforward approach, it still struggles to generate complex characters such as Korean, Japanese, and Chinese.
 GlyphByT5 series [19, 20] employ character-aware ByT5 encoder [17] and a new cross attention mechanism to compute the text region and image region separately.
-GlyphDraw [18] focuses on generating complex characters such as Chinese and proposes a sound baseline by introducing glyph condition for visual text generation.
-It uses the glyph image of the text and a location mask as conditions, allowing the model to control both the content and position of the generated text.
-However, it can generate only one line of text per image.
-TextDiffuser [5] trains a dedicated layout generation module to specifically generate character-level mask as the condition for the diffusion model, improving performance for Latin texts.
+Although replacing text encoders appears straightforward, it still struggles to generate complex characters such as Korean, Japanese, and Chinese.
+GlyphDraw [21] pioneered using glyph conditions and location masks for visual text generation of complex characters.
+It employs a glyph image and a location mask to control content and placement, though it remains limited to one line per image.
+TextDiffuser [6] introduces a dedicated layout generation module to produce character-level masks as diffusion model conditions, enhancing Latin text generation.
 But it still cannot generate non-Latin texts.
-GlyphControl [34] employs ControlNet [36] for visual text generation and uses a rendered glyph image as a control condition, making it possible to enable models like Stable Diffusion to generate texts and maintain their text-to-image capability at the same time.
-Nonetheless, GlyphControl can only generate texts in straight lines and often produces unwanted texts outside the text region.
-Following the above works, AnyText [32] proposes a robust and unified framework for visual text generation.
-It can generate high-quality multilingual texts in images, thanks to the text perceptual loss which takes the correctness of the generated text into account during training.
-However, AnyText performs poorly for generated text leaning more than $45$ degrees.
-TextGen [35] analyzes how the control information influences the generation in different timesteps.
-Diff-Text [37] uses off-the-shelf diffusion model and canny controlnet to generate visual texts, where the canny controlnet utilizes the canny image of text glyph to control visual text generation.
-Nonetheless, Diff-Text binds text position to predefined objects like ’sign’ or ’billboard’ in the prompt.
+TextDiffuser-2 [5] refines this approach by replacing character-level masks with bounding box coordinates, but this modification weakens its capacity for flexible visual text layout customization.
+GlyphControl [38] leverages ControlNet [40] and uses rendered glyph images as control conditions, enabling visual text generation in models like Stable Diffusion while preserving their core image synthesis capabilities.
+Nevertheless, GlyphControl is restricted to straight-line text layouts and frequently generates extraneous text artifacts.
+Building on these works, AnyText [36] proposes a unified framework for high-quality multilingual visual text generation.
+Its text perceptual loss directly evaluates text accuracy during training, improving correctness of generated visual text.
+However, AnyText struggles with rendering text rotated beyond 45 degrees.
+TextGen [39] examines how control signals affect image generation across timesteps.
+TextHarmony [42] unifies image comprehension, generation, and editing within a single framework by training a multimodal Large Language Model (LLM) to improve image understanding and generate more precise tokens for image synthesis.
+However, users face challenges in defining text layout during image generation, as current methods rely solely on textual prompts.
+In terms of text editing, its performance in complex backgrounds is unsatisfactory.
+Diff-Text leverages Canny ControlNet for glyph-based generation but ties text placement to predefined objects in prompts.
 It may fail if a prompt does not include predefined objects, leading to unnatural images.
@@
 Given user input on the leftmost side, which contains a prompt and a mask $l_{p}$ specifying positions for generating visual texts, we first split the $l_{p}$ using Divide and Conquer Strategy and obtain glyph image $l_{g}$ and flat position mask $\tilde{l}_{p}$.
-Then $\tilde{l}_{p}$ and $l_{g}$ are input to Semantic Rectification Branch and Structure Injection Branch respectively.
-In Semantic Rectification Branch, based on $\tilde{l}_{p}$, we render flat glyph $\tilde{l}_{g}$ and use them along with prompt and random noise $z_{T}$ to generate the latent with flat visual text.
+Then $\tilde{l}_{p}$ and $l_{g}$ are input to the Semantic Rectification Branch and Structure Injection Branch respectively.
+In the Semantic Rectification Branch, based on $\tilde{l}_{p}$, we render flat glyph $\tilde{l}_{g}$ and use them along with prompt and random noise $z_{T}$ to generate the latent with flat visual text.
 This latent serves as a semantic prior, providing rich semantic information for both the generation of the text and its background.
 $l_{g}$, on the other hand, is converted into the latent space as a structural prior for structural refinement of the text.
-Finally, the two prior combined together to guide the generation of the visual text in $l_{p}$.
+Finally, the two prior combined to guide the generation of the visual text in $l_{p}$.
 Given a prompt $y$ describing the image and a challenging position mask $l_{p}$, our goal is to generate an image based on $y$ that incorporates visual text at the specified positions dictated by $l_{p}$.
-Current visual text generation models struggle to generate visual texts in more challenging $l_{p}$ settings, such as tilted or curved texts layouts.
+Current visual text generation models struggle to generate visual texts in more challenging $l_{p}$ settings, such as tilted or curved text layouts.
 Our method aims to solve the problem based on two key insights: First, when the position mask $l_{p}$ is flat, current models can generate visual texts with high accuracy.
@@
 As shown in Fig.
-3, the semantic rectification branch first takes the masks $\tilde{l}_{p}$ reconfigured from $l_{p}$, prompt $y$, and random noise $z_{T}$ to generate the latent representation $z_{0}^{f}$, which contains reconfigured and flat visual text as a semantic prior.
+4, the semantic rectification branch first takes the masks $\tilde{l}_{p}$ reconfigured from $l_{p}$, prompt $y$, and random noise $z_{T}$ to generate the latent representation $z_{0}^{f}$, which contains reconfigured and flat visual text as a semantic prior.
 The structure injection branch then takes the rendered glyph image $l_{g}$ and $z_{t}$ to generate a structural prior.
@@
 3.1 Semantic Rectification Branch
-In the initial stages of the DDIM denoising process in the VGTM, clear text is created when applying a flat position mask, while blurry and unclear text is generated when using a slanted mask.
+In the initial stages of the DDIM denoising process in the VGTM, clear text is generated.
+This text remains accurate when a flat position mask is applied.
+However, when using a slanted mask, the text gradually becomes distorted.
 The reason for the distortion lies in the semantic drift in the text region when tilted.
-Motivated by this observation, we inherit the high quality flat text generation capabilities of the existing visual text model, adopting the latent of flat text as a constant reference to rectify the semantic information for challenging text generation.
+Motivated by this observation, we inherit the high-quality flat text generation capabilities of the existing visual text model, adopting the latent of flat text as a constant reference to rectify the semantic information for challenging text generation.
 Reference Branch for Semantic Rectification.
@@
 As shown in Fig.
-3, the flat position $\tilde{l}_{p}$ and corresponding glyph image $\tilde{l}_{g}$ are fed into the VTGM along with prompt $y$ and random noise $z_{T}$.
+4, the flat position $\tilde{l}_{p}$ and corresponding glyph image $\tilde{l}_{g}$ are fed into the VTGM along with prompt $y$ and random noise $z_{T}$.
 After $T$ denoising steps, we obtain the reference latent $z^{f}_{0}$.
@@
 $\tilde{z}_{t}=z_{f}\odot l_{p}+z_{t}\odot(1-l_{p}),$ (1)
-where $l_{p}$ is the position mask input by user.
+where $l_{p}$ is the position mask input by the user.
 This operation effectively rectifies the visual text using the accurate text in the $z_{f}$.
@@
 [22] demonstrated that AdaIN can be applied directly to Diffusion models for style transfer without any additional training.
-In our method, we also incorporate AdaIN, but with the goal of preserving image integrity against disruptions in latent distribution caused by the replacement operation:
+In our method, we also incorporate AdaIN, but to preserve image integrity against disruptions in latent distribution caused by the replacement operation:
 $\tilde{z}_{t}$ $={\rm AdaIN}(z_{f},z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$ (2)
@@
 To deal with the problem, we propose a divide-and-conquer strategy to effectively segment complex position masks into more manageable straight sections as shown in Fig.
-3.
+4.
 We use Bézier curves to define the upper and lower boundaries of $l_{p}$ and establish a baseline for the texts by averaging these curves.
@@
 With the assistance of this branch, we can generate highly accurate texts and create harmonious, diverse images by leveraging rich semantic information in both the visual texts and background.
-1 Input: A text prompt $y$, an input position image $l_{p}$, a flat reference position image $\tilde{l}_{p}$ and a set of iterations for refinement $\left\{t_{1},\ldots,t_{k}\right\}$ and a trained Visual Text Generation Model $\mathcal{M}$
-2 Output: Denoised vector $z_{0}$
-3   Render the glyph image $l_{g}$ and compute its latent $z_{g}$;
-4 Compute the reference latent $z^{f}_{0}$;
-5 Rotate and crop $z^{f}_{0}$ to get $z_{f}$ according to the transformation from $\tilde{l}_{p}$ to $l_{p}$;
-6 for $t=T,T-1,\ldots,1$ do
-7       if $t\in\left\{t_{1},\ldots,t_{k}\right\}$ then
-8             $\tilde{z}_{t}\leftarrow{\rm AdaIN}(z_{f},z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$;
-9             $\hat{z}_{t}\leftarrow\rho{\rm AdaIN}(z_{g},z_{t})+(1-\rho)\tilde{z}_{t}$;
-10             $\ddot{z}_{t}\leftarrow(\kappa_{t}\lambda\hat{z}_{t}+(1-\kappa_{t})z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$;
-11             Set $z_{t}\leftarrow\ddot{z}_{t}$;
-12
-13       end if
-14      Set $z_{t-1}\leftarrow\mathcal{M}(z_{t},y,t,l_{p},l_{g})$;
-15
-16 end for
-17
-18Return $z_{0}$
-Algorithm 1 STGen
 3.2 Structure Injection Branch
@@
 As shown in Fig.
-3, the structure injection branch feeds the rendered glyph image $l_{g}$ to Variational Autoencoder (VAE) [11] encoder to obtain the latent $z_{g}$.
+4, the structure injection branch feeds the rendered glyph image $l_{g}$ to Variational Autoencoder (VAE) [14] encoder to obtain the latent $z_{g}$.
 As an extension of semantic prior, we incorporate the $z_{g}$ as a structure prior into the latent $z_{t}$, using the same AdaIN operation above to regulate the range, which is represented as:
@@
 $\ddot{z}_{t}=(\kappa_{t}\lambda\hat{z}_{t}+(1-\kappa_{t})z_{t})\odot l_{p}+z_{t}\odot(1-l_{p}),$ (7)
-where $\lambda$ is a hyper-parameter and the $\kappa_{t}$ is a temporal factor that decays over time with each timestep.
-Together, they control the injection strength of the merged prior.
-The complete algorithm is represented as Algorithm 1.
+where $\lambda$ is a hyper-parameter and the $\kappa_{t}$ is a temporal factor that decays over time with each timestep. Together, they control the injection strength of the merged prior.
 4 Experiments
@@
 English TextDiffuser [6] 49.88 20.10 0.262 29.06 70.86 41.25 4.47 45.78 0.3091
+TextDiffuser-2 [5] 0.59 0.00 0.15 0.32 3.23 0.60 0.50 1.81 0.2989
+SD1.5+TextHarmony [42] 1.14 0.00 0.15 0.58 16.85 2.51 1.12 8.92 0.3090
 GlyphControl [38] 19.00 1.63 0.22 9.47 41.74 9.17 2.98 22.95 0.3206
@@
 Diff-Text [41] 40.13 26.42 8.85 28.38 61.44 45.37 17.94 45.91 0.2962
+Diff-Text+Ours 62.23 50.99 27.28 50.21 76.49 64.34 37.49 63.18 0.3018
 GlyphControl+Ours 39.19 2.63 1.05 19.48 55.57 11.83 6.74 31.17 0.3175
@@
 Chinese TextDiffuser [6] 5.41 4.07 0.09 4.07 56.81 49.37 43.01 52.47 0.3066
+TextDiffuser-2 [5] 0.06 0.08 0.00 0.05 7.28 4.81 6.04 6.49 0.2984
+SD1.5+TextHarmony [42] 0.00 0.00 0.00 0.00 14.35 8.50 12.00 12.60 0.3104
 GlyphControl [38] 2.33 0.24 0.0 1.41 50.98 15.58 35.94 40.24 0.3192
@@
 Diff-Text [41] 23.46 17.26 7.35 18.95 68.27 62.71 48.65 63.21 0.2950
+Diff-Text+Ours 42.13 38.43 22.42 37.47 81.15 71.84 59.69 74.91 0.2978
 GlyphControl+Ours 6.27 0.57 0.74 3.93 62.50 18.08 45.58 49.40 0.3168
@@
 Table 2: Quantitative Comparison between STGen and other competitors on both English and Chinese sets.
-All competitors are evaluated based on their officially released models.
+All competitors are evaluated based on their officially released code and models.
 Numbers in bold indicate the best performance, and underscored numbers indicate the second best.
-Baseline Baseline Preference Our Method Preference
-TextDiffuser 23.07% 76.93%
-GlyphControl 21.96% 78.04%
-SD1.5+SceneVTG 18.99% 81.01%
-AnyText 25.31% 74.69%
-Diff-Text 17.00% 83.00%
-Table 3: User study results.
-The participants were asked to select the best results based on image quality, the accuracy of the generated visual texts, and the similarity between the prompt and the generated images.
+indicates we adapt the method which is originally focused on other tasks into image generation.
+Method SEN.ACC $\uparrow$ NED $\uparrow$ FID $\downarrow$
+GlyphControl 37.10/3.27 66.80/8.45 37.84/34.36
+w/ STGen 70.92/11.81 84.68/24.72 32.00/32.99
+Diff-Text 56.11/29.49 75.28/48.67 69.24/62.61
+w/ STGen 68.55/39.58 81.55/55.61 66.77/61.59
+AnyText 72.39/69.23 87.67/83.96 33.54/31.58
+w/ STGen 75.53/70.00 89.38/84.57 33.83/32.17
+Table 3: Evaluation on vanilla AnyText Benchmark. In each cell of the table, the numbers on the right are results from the English set while the numbers on the right are from the Chinese set.
 4.1 Implementation Details
-Our framework is based on the pretrained AnyText model.
-During inference, we utilize a single RTX 3090 GPU for images in size of (512, 512).
-The framework achieves optimal performance when $\lambda$ is within the range $[-0.5,0.5]$ and $\rho$ is within $(0,2]$.
-We set $\kappa_{t}=10^{t-T}$ for effective guidance and harmonious text region boundary.
-Specifically, when $\lambda$ is in $[-0.5,0)$, the diversity of generated visual texts increases, while $\lambda$ in $(0,0.5]$ increases the variety of the image background.
-As shown in Tab.
-1, our method demonstrates robustness with respect to hyper-parameter sensitivity.
-In subsequent experiments, we set $\lambda$ to $0.5$ and $\rho$ to $0.5$ which has a balanced performance on both Chinese and English for evaluation.
+We use a single RTX 3090 GPU for images of size (512, 512).
+Optimal performance is achieved when $\lambda\in[-0.5,0.5]$ and $\rho\in(0,2]$.
+We set $\kappa_{t}=10^{t-T}$ for effective guidance and harmonious text region boundary.As shown in Tab.
+1, our method demonstrates robustness for hyper-parameter sensitivity.
+In subsequent experiments, we set $\lambda$ to $0.5$ and $\rho$ to $0.5$, which has a balanced performance on both Chinese and English for evaluation.
 4.2 Evaluation Setup
-Since there are currently no openly available datasets that focus on challenging visual texts generation, we establish a new benchmark for evaluation based on the AnyText-benchmark [32].
-For each text position mask, we randomly rotate the masks from the original benchmark and resolve any intersections, resulting in 984 prompts for LAION-word (English evaluation) and 919 prompts for Wukong-word (Chinese evaluation), each with corresponding position masks.
-Each prompt includes at least one visual text line and up to five visual text lines.
-During evaluation, we categorize the masks into three levels based on their angles: ‘easy’ (0°–30°), ‘medium’ (30°–60°), and ‘hard’ (60°–90°), facilitating multi-level assessments.
+Due to the lack of publicly available datasets focused on challenging visual text generation, we propose a new benchmark derived from the AnyText-benchmark [36].
+For each text position mask, we randomly rotate the original benchmark’s masks and resolve overlaps, yielding 984 prompts for LAION-word (English evaluation) and 919 prompts for Wukong-word (Chinese evaluation).
+During evaluation, masks are categorized into three difficulty levels based on rotation angles: easy (0°–30°), medium (30°–60°), and hard (60°–90°), enabling multi-level assessment of performance.
 Textual accuracy and background-text coherence are two main factors that determine the quality of slanted text generation, which we quantitatively evaluate through OCR accuracy.
 Following AnyText [36], we select the following two metrics for comparing OCR accuracy at word-level and character-level, respectively: (1) Sentence Accuracy (Sen.Acc); (2) Normalized Edit Distance (NED).
-We evaluated existing competing methods, including TextDiffuser [5], GlyphControl [34], SceneVTG [38], AnyText [32] and Diff-Text [37] using the benchmark and metrics mentioned above.
-It is worth noticing that the SceneVTG [38] cannot generate image from scratch.
-Therefore, we use the images generated by Stable Diffusion [27] with the same prompt as background image.
-For all the methods in the evaluation, we use the preprocessed mask using the Divide and Conquer strategy mentioned above to generate images to ensure fairness.
-Figure 4: Qualitative comparison of our method and state of the art models in both English and Chinese text generation.
-4.3 Qualitative Comparisons
-As presented in Fig.
-4, TextDiffuser tends to generate low-quality images, specifically in the second and the fifth columns.
-GlyphControl often produces texts outside the given text position, as can be seen in the fourth column.
-SceneVTG [38] exhibits severe text distortion, particularly in the first and third columns where the texts are extremely distorted and unrecognizable.
-Typical background occlusion is also observed in the second, third, and fourth columns, where the texts “Burning”, “Mamba” and “Twinkle” are obscured by non-textual elements such as lava, a human figure, and stars, respectively.
-This indicates that SceneVTG can only render texts against simple and clear backgrounds and does not integrate well with other text-to-image models.
-For Diff-text [37], there is a noticeable lack of coherence in the images due to insufficient consideration for harmony between the visual texts and the background, as particularly evidenced in the last 2 columns where the texts are rendered outside the bow and the watch.
-In AnyText [32], the generated visual texts tend to be distorted, this was caused by the loss of semantic and structure information during the inference.
-Moreover, in the first column, the text “Track” occludes with the train head, affecting both the text and the train.
-For our method, we can generate images with superior quality with accurate English and Chinese visual texts integrated even in complex scenarios like in the fourth and fifth columns.
-It’s worth noting that except for resolving text distortion, our method also improves the background occlusion between visual text and non-textual elements.
-For instance, in the third column of Fig.
-4, the text “Mamba” is originally disrupted by the human head using AnyText [32], while our method pushes the human head aside to generate a clear text and reasonable non-textual element.
+We evaluated existing competing methods, including TextDiffuser [6], TextDiffuser-2 [5], TextHarmony [42], GlyphControl [38], SceneVTG [43], AnyText [36] and Diff-Text [41] using the benchmark and metrics mentioned above.
+Notably, SceneVTG [43] cannot generate images from scratch.
+To address this, we use background images produced by Stable Diffusion [30] with identical prompts.
+For TextHarmony [42], its image generation model lacks support for positioning visual text at specific locations.
+Following its evaluation protocol on the AnyText Benchmark, we employ its visual text editing mode, which generates text on images where target positions are masked in black.
+For all the baselines in the evaluation, we use their officially released code and checkpoints.
+To ensure fairness, we preprocess the masks using the Divide and Conquer strategy mentioned above before generating images.
+Baseline Baseline Preference Ours Preference
+TextDiffuser [6] 23.07% 76.93%
+TextDiffuser-2 [5] 20.15% 79.85%
+GlyphControl [38] 21.96% 78.04%
+SceneVTG [43] 18.99% 81.01%
+TextHarmony [42] 18.75% 81.25%
+AnyText [36] 25.31% 74.69%
+Diff-Text [41] 17.00% 83.00%
+Table 4: User study results. Participants were asked to choose the best results based on image quality, accuracy of generated text within images, and prompt-image similarity.
+Figure 5: Qualitative comparison of our method and state-of-the-art models in both English and Chinese text generation.
 4.3 Quantitative Analysis
@@
 2, our method shows substantial improvements over the competitive methods in both English and Chinese across all levels.
-Notably, our method doesn’t require additional training and enhances the performance of GlyphControl and AnyText at the easy level.
+Notably, our method doesn’t require additional training and enhances the performance of Diff-Text, GlyphControl, and AnyText at the easy level.
 At the hard level, AnyText with our method yields results closest to the ground truth in the NED metric, with an improvement of approximately $10\%.$ over the baseline AnyText in both languages.
+As similarly shown in Tab.
+3, our methods outperform baselines on vanilla AnyText Benchmark except FID.
+Meanwhile, Diff-Text with our method shows an impressive $20\%$ improvement at the hard level in the English set.
 Text-Image Similarities.
@@
 2, our method increases the accuracy of visual text generation without compromising the baseline performance.
-Additionally, our method improves the text-image similarity slightly, due to our design that minimizes conflicts between the image composition and the visual text.
+Additionally, our method slightly improves text-image similarity due to our design, which minimizes conflicts between the image composition and the visual text.
 This leads to a more coherent layout for both the image and the text.
 User Study.
-Following [2], we conduct user study to comprehensively compare the generation results in Tab.
-3.
-We create 27 input sets with varying levels of difficulty in visual text generation.
-For each input set, we provide the respondents with our input prompt and position mask and two image alternatives: our result and a baseline (in random order).
-We ask them to select which image has the best quality and has the most accurate visual text.
-The final score is calculated based on the times the respondent choose the approach averaged on the total prompts.
-We collect 126 user judgments from a diverse group composed of both experts in the field and individuals with no specific background and report the percentage of votes.
+Following [2], we conduct a user study to comprehensively compare the generation results in Tab.
+4.
+We create 27 input sets of varying difficulty.
+For each set, participants receive our input prompt, position mask, and two images (our result and a baseline) in random order, and select the image with the best quality and most accurate visual text.
+The final score is the average number of selections per prompt.
+We gather 168 judgments from a diverse group of experts and non-experts and report vote percentages.
 As shown in Tab.
-3, our method is the most favored among all the circumstances.
-5 Ablation Study
+4, our method is preferred in all cases.
+Figure 6: More results on Background coherence and complex text layouts.
+The left two columns show how our approach seamlessly blends text with the background, while the remaining images highlight its ability to produce multi-sentence and circular text.
+4.4 Qualitative Comparisons
+As presented in Fig.
+5, TextDiffuser [6] tends to generate low-quality images, specifically in the second and the fifth rows.
+Textdiffuser-2 [5] fails to consistently generate legible text within specified bounding boxes.
+We suspect this occurs due to interference between textual and non-textual elements when using overly large bounding boxes.
+GlyphControl [38] frequently produces texts outside its designated area, particularly when the text is near the image boundary, as shown in the second and fourth row.
+SceneVTG [43] suffers from severe text distortion, especially in the first and third rows.
+Background occlusion is also evident in the second, third, and fourth rows, where the texts “Burning,” “Mamba,” and “Twinkle” are obscured by lava, a human figure, and stars, respectively.
+This suggests SceneVTG struggles with complex backgrounds and lacks seamless integration with other text-to-image models.
+TextHarmony [42] faces similar issues.
+In image editing mode, it struggles to properly fill blanks with complex shapes, leading to distorted text, as seen in the second row.
+For Diff-text [41], there is a noticeable lack of coherence in the images due to insufficient consideration for harmony between the visual texts and the background.
+AnyText [36] tends to produce distorted visual texts due to a loss of semantic and structural information during inference.
+For example, in the first row, the text “Track” overlaps with the train head, compromising both elements.
+In contrast, our method generates high-quality images with accurately rendered English and Chinese texts—even in challenging scenarios such as those in the fourth and fifth rows and Fig.
+6.
+Moreover, our approach not only corrects text distortion but also reduces background occlusion between textual and non-textual elements.
+For instance, in the third row of Fig.
+5, AnyText [36] disrupts the text “Mamba” with a human head, whereas our method repositions the head to produce clear text alongside well-integrated non-text elements.
 Figure 7: Ablation Study Visualization.
 The first column shows accuracy improvement with SRB.
-The second column demonstrates that while SIB enhances accuracy, it compromises coherence.
-The third column highlights that combining both branches yields accurate, coherent visual texts.
-Finally, AdaIN combination in the fourth column further refines image quality.
+The second reveals that SIB enhances accuracy but compromises coherence.
+The third confirms that combining both branches enhances both.
+Finally, the fourth shows that adding AdaIN further improves image quality.
+5 Ablation Study
 In this section, we demonstrate the effectiveness of our components through an ablation study on the medium level.
@@
 4.1.
-The analysis of each component will be presented in the following parts of this section.
 SRB without AdaIN.
 As shown in the first two rows of Tab.
-4, the addition of Semantic Rectification Branch improves the visual text accuracy.
+5, the addition of Semantic Rectification Branch improves the visual text accuracy.
 Similarly, as demonstrated by the first column of Fig.
-5, the generated visual text has a clear structure.
+7, the generated visual text has a clear structure.
 However, this also negatively impacts text-image consistency, as indicated by the slight drop in the CLIP score.
@@
 37.25 64.42 0.3027
-Table 4: Ablation study. The results validate the effectiveness of each component.
+Table 5: Ablation study.
 SIB without AdaIN.
 As illustrated in the second and third rows of Tab.
-4, Structure Injection Branch further improves the accuracy.
+5, Structure Injection Branch further improves the accuracy.
 As similarly shown in the third column of Fig.
-5, this branch can structurally improve the text structure, resulting in impressive improvement.
+7, this branch can structurally improve the text structure, resulting in impressive improvement.
 However, when applied alone, it may disrupt image coherence, as shown in the second column and third row of Fig.
-5, where the word ’Fun’ occludes the eye.
+7, where the word ’Fun’ occludes the eye.
 This issue arises from the lack of semantic information in the structural prior.
@@
 6 Conclusion
-In this work, we push the boundary of visual text generation by tackling the challenge of creating complex visual texts.
-We present a dual-branch method STGen, which decomposes the visual text generation into 2 branches: Semantic and Structure.
-In Semantic Rectification Branch, our method leverages the rich semantic information composed in the latent when generating accurate visual texts in simple scenarios to guide the generation in complex scenarios.
-In the Structure Injection Branch, we polish the visual text structure using the structural information extracted from glyphs at the same time.
-For challenging scenarios, we break them down into simpler steps to manage effectively.
-This approach can be integrated seamlessly with the existing visual text generation model using a dedicated ControlNet for generation.
-To exhibit its superior performance, we conduct extensive experiments on our proposed benchmark, demonstrating its excellence.
-Our approach offers a straightforward yet effective way to enhance the capabilities of existing visual text generation models.
-We believe STGen sets the stage for real-world applications, enabling more accurate and sophisticated visual text generation.
-Supplementary Material
-Figure 6: Comparison of predicted $x_{0}$ under different inference steps in various masks.
-The first column is the AnyText intermediate results under a flat mask, AnyText can predict the structure of the texts at the very early stage of inference.
-The rest of the columns show intermediates in a slanted mask using AnyText and other operations using different methods.
-Our method can revise the originally wrong texts and harmonize the background.
-7 Additional Background
-Latent Diffusion Models.
-The Latent Diffusion Models (LDM) denoise a noisy vector $z_{t}$ to $z_{0}$ that is mapped to an image $x_{0}$ through an autoencoder, based on a text prompt $y$.
-To sequentially remove the noise $\varepsilon$, a network $\varepsilon_{\theta}$ is trained to minimize the loss:
-$\mathcal{L}=\mathbb{E}_{z_{0},y,\varepsilon\sim\mathcal{N}(0,1),t}||\varepsilon-\varepsilon_{\theta}(z_{t},t,c(y))||_{2}^{2},$ (8)
-where the $c(y)$ is the conditioning embedding of the prompt $y$, $z_{t}$ is a noisy vector obtained by adding noise to $z_{0}$ according to the timestep $t$.
-During inference, provided a random noise vector $z_{T}$, the trained network iteratively removes the predicted noise to produce a latent $z_{0}$ for $T$ steps.
-Namely, we employ popular DDIM sampling [29] for inference:
-$z_{t-1}=$ $\sqrt{\alpha_{t-1}}\underbrace{(\frac{z_{t}-\sqrt{1-\alpha_{t}\epsilon_{\theta}(z_{t})}}{\sqrt{\alpha_{t}}})}_{\text{predicted $z_{0}$}}$ (9)
-$+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}\cdot\epsilon_{\theta}(z_{t})}.$
-The DDIM Sampling in general combines the predicted $z_{0}$ and $z_{t}$ to obtain the $z_{t-1}$.
-We mainly analyze the predicted $x_{0}$ that is decoded from predicted $z_{0}$ in the rest of the section.
-Adaptive Instance Normalization.
-The Adaptive Instance Normalization (AdaIN) [10] is previously adopted in the task of style transfer.
-It substitutes mean and standard deviation of the activations from each CNN [13] filter of the original image with those of the style image as follows:
-${\rm AdaIN}(x,y)=\sigma(y)(\frac{x-\mu(x)}{\sigma(x)})+\mu(y),$ (10)
-where $x$ and $y$ stand for the activation of a CNN filter from the content image and style image, $\mu(\cdot)$ and $\sigma(\cdot)$ are channel-wise mean and standard deviation.
-Masui et al.
-[19] further prove that the AdaIN can be directly applied to the LDM without any additional training for style transfer.
-8 Analysis of Predicted $x_{0}$
-As shown in Fig.
-6, AnyText can produce clear and intact visual texts in structure even in the first inference step.
-However, once tilted, the prediction becomes blurry and wrong as in the second column.
-The following columns show the effectiveness of our method.
-SRB can improve the structure.
-The SIB further improves the structure but produces visual texts that are not coherent with the background.
-Figure 7: Results under different lambda.
-When $\lambda$ is -0.5, we obtain colorful and styled visual texts in the first column.
-When $\lambda$ is 0.5, we produce images with rich background.
-When $\lambda$ is 0, the method degrades to vanilla AnyText.
-Language $\lambda$ $\rho$ Sen.ACC NED CLIP score
-easy medium hard total easy medium hard total
-English -0.5 0.5 71.33 35.51 6.14 44.88 87.23 62.98 22.79 64.11 0.3005
-0.5 0.5 71.25 37.25 6.60 45.43 87.54 64.42 24.56 65.10 0.3027
-0.5 0.25 71.37 37.11 5.32 45.12 87.10 63.80 21.19 63.85 0.3036
-0.5 0.75 69.60 38.39 7.16 45.10 87.33 65.54 25.94 65.65 0.3006
-0.5 1.25 61.48 35.16 6.33 40.18 85.38 63.73 25.44 64.11 0.3008
-0.5 1.5 69.47 38.92 6.48 45.01 87.23 65.03 25.43 65.34 0.3009
-0.5 1.75 69.19 38.17 6.56 44.70 87.18 65.31 25.40 65.39 0.3011
-0.5 2.0 68.88 37.96 6.71 44.53 87.07 64.85 25.26 65.17 0.3009
-Chinese -0.5 0.5 68.61 35.02 7.54 49.28 95.17 85.23 66.05 87.29 0.3067
-0.5 0.5 69.22 35.58 8.55 49.96 95.37 85.85 68.77 88.08 0.3071
-0.5 0.25 71.35 35.42 6.16 50.70 95.46 85.88 67.37 87.86 0.3076
-0.5 0.75 66.12 37.79 10.66 49.05 94.54 85.77 70.38 87.89 0.3058
-0.5 1.25 64.28 38.11 10.57 48.03 94.31 86.14 71.58 88.07 0.3060
-0.5 1.5 63.85 37.05 9.83 47.40 94.24 85.52 70.26 87.64 0.3061
-0.5 1.75 65.53 38.11 10.29 48.12 94.29 85.85 70.21 87.73 0.3064
-0.5 2.0 65.02 35.91 10.11 47.88 94.44 85.37 69.78 87.63 0.3059
-Table 5: Detailed sensitivity analysis for $[\lambda,\rho]$.
-9 Additional Results
-Here we provide additional results of our method. Our code will be released upon acceptance.
-9.1 Additional Quantitative Results
-Hyper-parameters Sensitivity
-As shown in the Tab.
-5, our method shows robustness in hyper-parameters change.
-The best performance is obtained when $\rho$ is 0.75 at the hard level.
-But the method achieves the best performance in the easy level of Chinese and English set when $\rho$ is 0.25.
-To ensure fairness, we choose the performance when $\rho$ is 0.5 which has a balanced performance for comparison.
-9.2 Additional Qualitative Results
-Choice of $\lambda$
-We can observe in Fig.
-7 that when $\lambda$ is in the range of $(0,0.5]$ we can achieve background with more diversity and in the range of $[-0.5,0)$ the diversity of texts is increased consistent with what we mentioned in Sec.
-4.1.
-Coherence with Background
-As shown in Fig.
-8, thanks to the rich semantic information contained in the latent with flat visual texts, we can generate visual texts that are perfectly blended in with the background.
-The text ‘Harvest’ in the bottom right is more harmonious with the background and is closest to the description “interwoven”.
-Visual Text Diversity
-As shown in the first row of Fig.
-10, with proper control, our method can generate texts in various styles and colors.
-The second row shows that the generated visual texts remain to be accurate under 2 different image styles.
-The third row demonstrates the robustness of our method under different challenging masks.
-The fourth row shows our ability to generate various visual texts using the same prompt.
-Additional Comparison with baseline.
-We present more qualitative comparisons on simple slanted or curved situations in Fig.
-11 and Fig.
-12, and complex scenarios in Fig.
-13.
-Our method can generate high-quality images with accurate visual texts in all the scenarios.
-10 Details of User Study
-An illustration of our user study is provided in Fig.
-9.
-Since the caption in the English (LAION) set are often hard for users to understand (e.g.
-“carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel”).
-To produce meaningful prompts that can help the user understand, we employ LLMs to generate image prompts for user study.
-Here we provide a simplified template of our instructions and examples:
-1.Task instruction
-I am working on a text-to-image generation task. Each Prompt should include a scene description and contain one or two sets of words that need to be present in the image, enclosed in double quotes.
-2. Examples
-A raccoon stands in front of the blackboard with the words “Deep Learning” written on it …… 3.Trigger CoT reasoning ability of LLMs
-Reasoning: Let’s think step by step……
-11 Limitations
-Our method still inherits the problems of the baseline AnyText, such as the unsatisfactory ability to generate texts on a small scale or to generate texts in a specific font.
-Restricted by the latent size, our method may not acquire accurate texts when the text region is too small.
-Figure 8: Examples of texts blend perfectly with background.
-Visual text generated by AnyText in the first row is often blurry and wrong.
-While our method can not only generate accurate visual text but the visual text is also blended with the background perfectly.
-Figure 9: User study print screen.
-Figure 10: Diversity demonstration.The first row demonstrates various fonts and colors a single prompt can generate.
-The second row shows same texts with different prompts can still generate coherent and reasonable images.
-We use different masks in the third row.
-The fourth row illustrates the ability to simply change the language of the generated texts.
-Figure 11: Additional slanted results for our method.
-Figure 12: Additional curved results for our method.
-Figure 13: Additional complex results for our method.
+We advance visual text generation by tackling the challenge of complex text synthesis.
+Our proposed STGen introduces a dual-branch approach: the Semantic Rectification Branch, which refines text generation using latent extracted from simpler scenarios, and the Structure Injection Branch, which enhances text structure by incorporating latent of glyph image.
+For highly challenging cases, we break them into more manageable cases.
+Integrated via a dedicated ControlNet, STGen seamlessly enhances existing models.
+Extensive experiments on our benchmark confirm its superior performance, making STGen a promising step toward real-world applications.